# Estimators for multivariate allometric regression model

### v2

@@ 1 Introduction
The multivariate regression model, a regression model with multiple response variables and multiple explanatory variables, is one of the most popular models in multivariate analysis; see Anderson (2003), Johnson and Wichern (2007), Rencher and Christensen (2012), and so on. For the model, the method of least squares works well. When additional assumptions such as low rankness of the coefficient matrix exist, there is another estimator that is generally better than the ordinary least squares estimator. In particular, we consider the multivariate allometric regression model, proposed in Tarpey and Ivey (2006), which is a reduced rank regression model of rank one with the constraint that the differences of the mean vectors of response variables for different values of explanatory variables are in the direction of the first principal eigenvector of the covariance matrix. The first principal eigenvector is a key quantity to infer the multivariate allometric regression model, so we consider the estimation of this quantity. Specifically, we define a class of estimators that includes natural conventional estimators, evaluate the mean squared errors of the estimators contained in this class, compare them from the viewpoint of the consistency, propose a new estimator, and examine the proposed estimator through numerical simulations.
@@
$\bm{Y}=\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{B}+\bm{E}\ \ \ {\rm with}\ \mathsf{E}[{\rm vec}(\bm{E})]=\bm{0}_{np},\ \mathsf{V}[{\rm vec}(\bm{E})]=\bm{\Sigma}\otimes\bm{I}_{n}$ ( 1.1)
@@
$\bm{y}_{i}=\bm{\mu}+\bm{B}^{\prime}\bm{x}_{i}+\bm{e}_{i}\ \ \ {\rm with}\ \mathsf{E}[\bm{e}_{i}]=\bm{0}_{p},\ \mathsf{V}[\bm{e}_{i}]=\bm{\Sigma},\ \ \ i=1,2,\dots,n.$
@@
$\bm{Y}=\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\ \ \ {\rm with}\ \mathsf{E}[{\rm vec}(\bm{E})]=\bm{0}_{np},\ \mathsf{V}[{\rm vec}(\bm{E})]=\bm{\Sigma}\otimes\bm{I}_{n}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}\otimes\bm{I}_{n},$ ( 1.2)
@@
The multivariate allometric regression model was proposed in Tarpey and Ivey (2006) as an extension of the allometric extension model (Flury, 1997; Hills, 2006) to the context of multivariate regression. The allometric extension model means that several population distributions share a first principal eigenvector of their covariance matrices and the differences of their mean vectors are in the direction of the first principal eigenvector, which has been applied to various areas including biometrics (Klingenberg and Froese, 1991; Flury, 1997). Bartoletti et al. (1999) proposed a hypothesis testing procedure for the allometric extension model based on a $\chi^{2}$ statistic in large sample situations. Kurata et al. (2008) derived conditions for multivariate conditional distributions to form the allometric extension model. Matsuura and Kurata (2014) discussed principal points of multivariate mixture distributions of the allometric extension model. Tsukuda and Matsuura (2023) proposed hypothesis testing procedures for the allometric extension model in high-dimensional and strongly spiked eigenvalue situations. Kawamoto et al. (2023) evaluated the performance of spectral clustering algorithm when it is used for a random sample from the mixture of two subpopulations that are allometric extensions of each other.
@@
This paper studies the estimation of $\bm{\gamma}_{1}$ in (1.2). From the viewpoint of principal component analysis, inferring the first principal eigenvector $\bm{\gamma}_{1}$ is of special interest. Moreover, although we focus on the estimation of $\bm{\gamma}_{1}$, precise estimation of $\bm{\gamma}_{1}$ may lead to precise estimation of $\bm{B}=\bm{\alpha}\bm{\gamma}_{1}^{\prime}$. In fact, once an estimator $\hat{\bm{\gamma}}_{1}$ of $\bm{\gamma}_{1}$ is obtained, an estimator of $\bm{B}$ is obtained as $\hat{\bm{B}}(\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime})=\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime}$ if $\bm{X}^{\prime}\bm{X}$ is invertible. In the reduced rank regression, such estimators are generally better than the ordinary least squares estimator $\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}$; see Anderson (1999). Also when (1.2) is supposed, $\hat{\bm{B}}(\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime})$ is considered to be better than the ordinary least squares estimator in general. This is another motivation to focus on estimating $\bm{\gamma}_{1}$. In particular, the consistency up to sign of $\hat{\bm{\gamma}}_{1}$ is important. That is because, e.g., under some adequate asymptotic regimes, if $\hat{\bm{\gamma}}_{1}$ is consistent, then $\hat{\bm{B}}(\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime})$ is also consistent.
In this paper, we deal with a class of estimators of $\bm{\gamma}_{1}$ based on weighted matrices of two independent sum-of-squares matrices: regression sum-of-squares matrix and residual sum-of-squares matrix. Non-asymptotic upper bounds are established for the mean squared errors of the estimators, and the weight value minimizing the upper bound for the mean squared error is derived. Subsequently, we discuss the consistency of the estimators under weak identifiability regimes under which $\lambda_{1}-\lambda_{2}\to 0$ as $n\to\infty$, and “large $p$, large $n$” regimes under which $p$ and $n$ tend to $\infty$ simultaneously. As remarked in Paindaveine et al. (2020), properties of principal component analysis under weak identifiability regimes have not been sufficiently investigated. Under large $p$, large $n$ regimes that include so-called high-dimensional asymptotic regimes $n,p\to\infty$ with $p/n\to\infty$, we consider weak identifiability regimes and a sort of spiked eigenvalue regimes simultaneously. In the context of multivariate regression, large $p$, large $n$ regimes have been discussed in a lot of works such as Bing and Wegkamp (2019), Bunea et al. (2011), Chen et al. (2013), Giraud (2011), She and Chen (2017), Yanagihara et al. (2015), Yuan et al. (2007), and so on. Furthermore, numerical results are presented to compare the proposed estimator with some conventional estimators.
@@ 2 Conventional estimators
Note that $\hat{\bm{\gamma}}_{1}^{(T)}$ is the first principal eigenvector in the usual principal component analysis for the multivariate location-scale model that is given by $\bm{\alpha}=\bm{0}_{q}$ in (1.2).
@@
In our setup, $\hat{\bm{\gamma}}_{1}^{(E)}$ is the first principal eigenvector of a Wishart matrix with $n-q-1$ degrees of freedom and scale matrix parameter $\bm{\Sigma}$. Thus, if $q$ is fixed, properties of $\hat{\bm{\gamma}}_{1}^{(E)}$ can be interpreted as general properties for the Wishart distribution.
@@ 3 Estimators based on weighted matrices of $\bm{S}_{R}$ and $\bm{S}_{E}$
$\mathsf{E}\left[\left\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right\|_{\rm F}^{2}\right]=\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+2(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}.$
@@
$=\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)^{\prime}\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)\right)\right]=\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)^{2}\right)\right]$
@@
First, we evaluate $\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\right)^{2}\right)\right]$ as
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\right)^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left\{\bm{Y}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}-\left(q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\right)\right\}^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left\{\left(\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\right)^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\left(\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\right)-\left(q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\right)\right\}^{2}\right)\right]$
@@
We note that $\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}$ is distributed as the $p$-dimensional Wishart distribution $\mathcal{W}_{p}\left(q,\bm{\Sigma}\right)$ with $q$ degrees of freedom and scale matrix parameter $\bm{\Sigma}$. It is well known that if $\bm{W}\sim\mathcal{W}_{p}\left(q,\bm{\Sigma}\right)$, then $\mathsf{E}\left[\bm{W}\right]=q\bm{\Sigma}$ and $\mathsf{E}\left[{\rm tr}\left(\bm{W}^{2}\right)\right]=(q+q^{2}){\rm tr}\left(\bm{\Sigma}^{2}\right)+q\left({\rm tr}\left(\bm{\Sigma}\right)\right)^{2}$. Moreover, as it will be seen in Lemma 3.2 below, it also holds that $\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]=\bm{0}_{p\times q}$. Hence, letting $\bm{W}=\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}$, we have
@@
$=\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\Bigl{(}\bm{W}^{2}-q\bm{W}\bm{\Sigma}-q\bm{\Sigma}\bm{W}+q^{2}\bm{\Sigma}^{2}$
$\quad+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\Bigr{)}\Bigr{)}\Bigr{]}$
@@
Next, we evaluate $\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}]\right)^{2}\right)\right]$. Since $\bm{S}_{E}\sim\mathcal{W}_{p}\left(n-1-q,\bm{\Sigma}\right)$, it holds that
@@
=$=\left\{(n-1-q)+(n-1-q)^{2}\right\}{\rm tr}\left(\bm{\Sigma}^{2}\right)+(n-1-q)\left({\rm tr}\left(\bm{\Sigma}\right)\right)^{2}-(n-1-q)^{2}{\rm tr}\left(\bm{\Sigma}^{2}\right)$
@@
=$=\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+2(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}.$
@@
Let $\bm{g}_{1}^{\prime},\bm{g}_{2}^{\prime},\dots,\bm{g}_{q}^{\prime}$ be the row vectors of $\bm{G}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1/2}\bm{X}^{\prime}\bm{E}$, that is, $\bm{G}=\left(\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}\right)^{\prime}$, where $(\bm{X}^{\prime}\bm{X})^{-1/2}$ is the $q\times q$ symmetric matrix satisfying that $\{(\bm{X}^{\prime}\bm{X})^{-1/2}\}^{2}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1}$. Then, from $\mathsf{V}[{\rm vec}(\bm{G})]=\bm{\Sigma}\otimes\bm{I}_{q}$, it follows that $\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}$ are independent and identically distributed as $\mathcal{N}_{p}(\bm{0}_{p},\bm{\Sigma})$, and thus we see that
$\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]=\mathsf{E}\left[\bm{G}^{\prime}\bm{G}\bm{G}^{\prime}\left(\bm{X}^{\prime}\bm{X}\right)^{1/2}\right]=\mathsf{E}\left[\sum_{i=1}^{q}\bm{g}_{i}\bm{g}_{i}^{\prime}\left(\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}\right)\right]\left(\bm{X}^{\prime}\bm{X}\right)^{1/2}=\bm{0}_{p\times q}.$
@@
Now, we establish an upper bound, without any unspecified constant, for the mean squared error (up to sign) of the estimator $\hat{\bm{\gamma}}_{1}(w)$:
@@
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}.$ ( 3.1)

### v4

@@ 1 Introduction
The multivariate regression model, a regression model with multiple response variables and multiple explanatory variables, is one of the most popular models in multivariate analysis; see Anderson (2003), Johnson and Wichern (2007), Rencher and Christensen (2012), and so on. For the model, the method of least squares works well. When additional assumptions such as low rankness of the regression coefficient matrix exist, there is another estimator that is generally better than the ordinary least squares estimator. In particular, we consider the multivariate allometric regression model, proposed in Tarpey and Ivey (2006), which is a reduced rank regression model of rank one with the constraint that the differences of the mean vectors of response variables for different values of explanatory variables are in the direction of the first principal eigenvector of the covariance matrix. The first principal eigenvector is a key quantity to infer the multivariate allometric regression model, so we consider the estimation of this quantity. Specifically, we define a class of estimators that includes natural conventional estimators, evaluate the mean squared errors of the estimators contained in this class, compare them from the viewpoint of the consistency, propose a new estimator, and examine the proposed estimator through numerical simulations.
@@
$\bm{Y}=\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{B}+\bm{E}\ \mbox{with}\ \mathsf{E}[{\rm vec}(\bm{E})]=\bm{0}_{np},\ \mathsf{V}[{\rm vec}(\bm{E})]=\bm{\Sigma}\otimes\bm{I}_{n}$ ( 1.1)
@@
$\bm{y}_{i}=\bm{\mu}+\bm{B}^{\prime}\bm{x}_{i}+\bm{e}_{i}\ \mbox{with}\ \mathsf{E}[\bm{e}_{i}]=\bm{0}_{p},\ \mathsf{V}[\bm{e}_{i}]=\bm{\Sigma},\ i=1,2,\dots,n.$
@@
$\bm{Y}=\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}$
$\ \mbox{with}\ \mathsf{E}[{\rm vec}(\bm{E})]=\bm{0}_{np},\ \mathsf{V}[{\rm vec}(\bm{E})]=\bm{\Sigma}\otimes\bm{I}_{n}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}\otimes\bm{I}_{n},$ ( 1.2)
@@
The multivariate allometric regression model was proposed as an extension of the allometric extension model (Flury, 1997; Hills, 2006) to the context of multivariate regression in Tarpey and Ivey (2006). The allometric extension model means that several population distributions share a first principal eigenvector of their covariance matrices and the differences of their mean vectors are in the direction of the first principal eigenvector, which has been applied to various areas including biometrics (Klingenberg and Froese, 1991; Flury, 1997). Bartoletti et al. (1999) proposed a hypothesis testing procedure for the allometric extension model based on a $\chi^{2}$ statistic in large sample situations. Kurata et al. (2008) derived conditions for multivariate conditional distributions to form the allometric extension model. Matsuura and Kurata (2014) discussed principal points of multivariate mixture distributions of the allometric extension model. Tsukuda and Matsuura (2023) proposed hypothesis testing procedures for the allometric extension model in high-dimensional and strongly spiked eigenvalue situations. Kawamoto et al. (2023) evaluated the performance of spectral clustering algorithm when it is used for a random sample from the mixture of two subpopulations that are allometric extensions of each other.
@@
This paper studies the estimation of $\bm{\gamma}_{1}$ in (1.2). From the viewpoint of principal component analysis, inferring the first principal eigenvector $\bm{\gamma}_{1}$ is of special interest. Moreover, although we focus on the estimation of $\bm{\gamma}_{1}$, precise estimation of $\bm{\gamma}_{1}$ may lead to precise estimation of $\bm{B}=\bm{\alpha}\bm{\gamma}_{1}^{\prime}=\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}$. In fact, once an estimator $\hat{\bm{\gamma}}_{1}$ of $\bm{\gamma}_{1}$ is obtained, an estimator of $\bm{B}$ is obtained as $\hat{\bm{B}}_{OLS}\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime}$ if $\bm{X}^{\prime}\bm{X}$ is invertible, where $\hat{\bm{B}}_{OLS}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}$ is the ordinary least squares estimator. In the reduced rank regression, such estimators are generally better than $\hat{\bm{B}}_{OLS}$; see Anderson (1999). Also when (1.2) is supposed, it seems natural to consider that $\hat{\bm{B}}_{OLS}\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime}$ is better than $\hat{\bm{B}}_{OLS}$ in general. This is another motivation to focus on estimating $\bm{\gamma}_{1}$. In particular, the consistency up to sign of $\hat{\bm{\gamma}}_{1}$ is important. That is because, e.g., under some adequate asymptotic regimes, if $\hat{\bm{\gamma}}_{1}$ is consistent, then $\hat{\bm{B}}_{OLS}\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime}$ is also consistent.
In this paper, we deal with a class of estimators of $\bm{\gamma}_{1}$ based on weighted matrices of two independent sum-of-squares matrices: regression sum-of-squares matrix and residual sum-of-squares matrix. Non-asymptotic upper bounds are established for the mean squared errors of the estimators, and the weight value minimizing the upper bound for the mean squared error is derived. Subsequently, we discuss the consistency of the estimators under weak identifiability regimes under which $\lambda_{1}-\lambda_{2}\to 0$ as $n\to\infty$, and “large $p$, large $n$” regimes under which $p$ and $n$ tend to $\infty$ simultaneously. As remarked in Paindaveine et al. (2020), properties of principal component analysis under weak identifiability regimes have not been sufficiently investigated. Under large $p$, large $n$ regimes that include so-called high-dimensional asymptotic regimes $n,p\to\infty$ with $p/n\to\infty$, we consider weak identifiability regimes and a sort of spiked eigenvalue regimes simultaneously. In the context of multivariate regression, large $p$, large $n$ regimes have been discussed in a lot of works such as Bing and Wegkamp (2019), Bunea et al. (2011), Chen et al. (2013), Giraud (2011), She and Chen (2017), Yanagihara et al. (2015), Yuan et al. (2007), and so on. Furthermore, numerical results are presented to compare the proposed estimator with some conventional estimators. Finally, a illustrative real data example is provided.
@@ 2 Conventional estimators
As stated above, $\mathrm{rank}(\bm{X})=q$ is assumed throughout the paper. When $\mathrm{rank}(\bm{X})<q$, we may extend analysis procedures studied in this paper by replacing $(\bm{X}^{\prime}\bm{X})^{-1}$ with the Moore–Penrose inverse of $\bm{X}^{\prime}\bm{X}$ or removing redundant explanatory variables with modifying the corresponding parts such as degrees of freedom.
Remark 2.
Note that $\hat{\bm{\gamma}}_{1}^{(T)}$ is the first principal eigenvector in the usual principal component analysis for the multivariate location-scale model with finite second moment that is given by $\bm{\alpha}=\bm{0}_{q}$ in (1.2).
@@
In our setup, $\hat{\bm{\gamma}}_{1}^{(E)}$ is the first principal eigenvector of a Wishart matrix with $n-1-q$ degrees of freedom and scale matrix parameter $\bm{\Sigma}$. Thus, if $q$ is fixed, properties of $\hat{\bm{\gamma}}_{1}^{(E)}$ can be interpreted as general properties for the Wishart distribution.
@@ 3 Estimators based on weighted matrices of $\bm{S}_{R}$ and $\bm{S}_{E}$
=$=\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}$
$\quad+2(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}.$
@@
$=\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)^{\prime}\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)^{2}\right)\right]$
@@
First, we evaluate $\mathsf{E}[{\rm tr}((\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])^{2})]$ as
$\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\bigl{(}\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\bigr{)}^{2}\Bigr{)}\Bigr{]}$
$=\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\bigl{\{}\bm{Y}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}-\bigl{(}q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bigr{)}\bigr{\}}^{2}\Bigr{)}\Bigr{]}$
$=\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\Bigl{\{}\left(\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\right)^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\left(\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\right)$
$\qquad-\left(q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\right)\Bigr{\}}^{2}\Bigr{)}\Bigr{]}$
@@
We note that $\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}$ is distributed as the $p$-dimensional Wishart distribution $\mathcal{W}_{p}\left(q,\bm{\Sigma}\right)$ with $q$ degrees of freedom and scale matrix parameter $\bm{\Sigma}$. It is well known that if $\bm{W}\sim\mathcal{W}_{p}\left(q,\bm{\Sigma}\right)$, then
$\mathsf{E}\left[\bm{W}\right]=q\bm{\Sigma},\quad\mathsf{E}\left[{\rm tr}\left(\bm{W}^{2}\right)\right]=(q+q^{2}){\rm tr}\left(\bm{\Sigma}^{2}\right)+q\left({\rm tr}\left(\bm{\Sigma}\right)\right)^{2}.$
Moreover, as it will be seen in Lemma 3.2 below, it also holds that
$\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]=\bm{0}_{p\times q},$
where $\bm{0}_{p\times q}$ is the $p\times q$ zero matrix. Hence, letting $\bm{W}=\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}$, we have
@@
$=\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\bm{W}^{2}-q\bm{W}\bm{\Sigma}-q\bm{\Sigma}\bm{W}+q^{2}\bm{\Sigma}^{2}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}$
$\qquad+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\Bigr{)}\Bigr{]}$
@@
Next, we evaluate $\mathsf{E}[{\rm tr}((\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])^{2})]$. Since $\bm{S}_{E}\sim\mathcal{W}_{p}\left(n-1-q,\bm{\Sigma}\right)$, it holds that
@@
=$=\left\{(n-1-q)+(n-1-q)^{2}\right\}{\rm tr}\left(\bm{\Sigma}^{2}\right)$
$\quad+(n-1-q)\left({\rm tr}\left(\bm{\Sigma}\right)\right)^{2}-(n-1-q)^{2}{\rm tr}\left(\bm{\Sigma}^{2}\right)$
@@
=$=\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}$
$\quad+2(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}.$
@@
Let $\bm{g}_{1}^{\prime},\bm{g}_{2}^{\prime},\dots,\bm{g}_{q}^{\prime}$ be the row vectors of $\bm{G}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1/2}\bm{X}^{\prime}\bm{E}$, that is, $\bm{G}=\left(\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}\right)^{\prime}$, where $(\bm{X}^{\prime}\bm{X})^{-1/2}$ is the $q\times q$ symmetric matrix satisfying that
$\{(\bm{X}^{\prime}\bm{X})^{-1/2}\}^{2}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1}.$
Then, from $\mathsf{V}[{\rm vec}(\bm{G})]=\bm{\Sigma}\otimes\bm{I}_{q}$, it follows that $\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}$ are independent and identically distributed as $\mathcal{N}_{p}(\bm{0}_{p},\bm{\Sigma})$, and thus we see that
$\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]$
$=\mathsf{E}\left[\bm{G}^{\prime}\bm{G}\bm{G}^{\prime}\left(\bm{X}^{\prime}\bm{X}\right)^{1/2}\right]=\mathsf{E}\left[\sum_{i=1}^{q}\bm{g}_{i}\bm{g}_{i}^{\prime}\left(\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}\right)\right]\left(\bm{X}^{\prime}\bm{X}\right)^{1/2}=\bm{0}_{p\times q}.$
@@
Now, we establish an upper bound without any unspecified constant for the mean squared error (up to sign) of the estimator $\hat{\bm{\gamma}}_{1}(w)$:
@@
$\leq\frac{8a\{q(1-2w)+(n-1)w^{2}\}+16bc(1-w)^{2}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}},$ ( 3.1)
where
$a={\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2},\ b=\lambda_{1}+{\rm tr}(\bm{\Sigma}),\ c=\|\bm{X}\bm{\alpha}\|^{2},\ d=\lambda_{1}-\lambda_{2}.$ ( 3.2)